You purchased 200 shares of CNX common stock on margin at $51 per share. Assume the initial margin is 50% and the maintenance margin is 34%. You will get a margin call if the stock drops below ________.

Assume the stock pays no dividends, and you pay no interest on the margin loan. PLEASE SHOW WORK
Answer #1

HI,

Lets say the price =P when we get a margin call

So your equity = 200* share price - 50%*200*51

=200P - 5100

margin = (200P-5100)/200P =0.34

200P -5100 = 68P

P =38.64

So we get a margin call if stock drops below $38.64

Thanks
Answer #2

ANSWER :


Per share basis calculations :


Initial margin = 50%


Stock price = $51


So, Initial margin amount = 51*0.50 = 25.50 ($)


Maintenance margin = 34% 


Let the stock price be P when margin call is effected.


So,


Maintenance margin amount = Initial margin amount - Loss

=> P * 0.34 = 25.50 - (51 - P)

=>  P -  0.34 P = 25.50

=>  0.66 P = 25.50

=> P = 25.50 / 0.66

=> P =  38.64  ($) 


Hence, margin call is received when stock price drops below $38.64 (ANSWER).


answered by: Tulsiram Garg